Two short magnets each of moment 10A-m square are placed in end-on position so that their centers are 0.1m apart.The force between them is

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
F=u0/4 pi (6M1M2/ X4)=10-76(10*10)/0.14=0.6N
 
In case of any difficulty please feel free to ask
Regards
Arun (askIITians forum expert)